OpenStudy (anonymous):

Determine the nature of the solutions of the equation x^2-12x+36=0

OpenStudy (jamesj):

Well, what is the value of b^2 - 4ac for this equation and what does that value imply?

hero (hero):

Cali, you don't need to delete your questions anymore.

OpenStudy (anonymous):

my questions are filling up the board and making it go slower right?

OpenStudy (anonymous):

i think you are supposed to recognize this as a "perfect square" i.e. \[(x-b)^2=x^2-12x+36=0\] see if you recognize b.

hero (hero):

I recognize b

OpenStudy (anonymous):

we know you do hero, your good at this crap. i am dumb as hell at math

hero (hero):

Well, you could figure it out by factoring

OpenStudy (anonymous):

and everyone who has been so kind in the past to me are now making fun of me for not understanding

hero (hero):

I hope I wasn't one of them. Why do you always think people are making fun of you. You shouldn't think like that. There are many people who don't understand this stuff.

OpenStudy (anonymous):

This equation either has one real, two real, or two non real solutions

OpenStudy (anonymous):

x=6

hero (hero):

Yes, it only has one solution. That is correct. Looks like James J had the right approach all along

OpenStudy (anonymous):

thanks, it took some time, but figured it out